A scale reads 274 N when a piece of brass is hanging from it. What does it read (in N) when it is lowered so that the brass is submerged in water?
1
Expert's answer
2019-04-15T09:52:44-0400

The weight of the piece of brass in air is "F_a = \rho_b V g", where "\rho_b" is the brass density, "V" is its volume, and "g" is the acceleration of gravity. By the Archimedes law, the weight of the same piece of brass submerged in water is "F_w = \rho_b V g - \rho_w V g", where "\rho_w" is the density of water. The last formula can be rewritten as

"F_w = \rho_b V g \left( 1 - \frac{\rho_w}{ \rho_b} \right) = F_a \left( 1 - \frac{\rho_w}{ \rho_b} \right) \, ."


Substituting here the corresponding densities "\rho_b \approx 8.5\, \text{g} / \text{cm}^3" and "\rho_w \approx 1\, \text{g} / \text{cm}^3", we obtain the weight of the brass in water: "F_w \approx 242\, \text{N}".


Answer: 242 N.